7) Suppose 5.2% of students enrolled in a local college do not graduate on time. Find the mean and standard deviation of 3500 students who do not graduate on time.
Answer by stanbon(75887)

Suppose 5.2% of students enrolled in a local college do not graduate on time. Find the mean and standard deviation of 3500 students who do not graduate on time.
------
Binomial Problems:
mean = np = 3500*0.052 = 182
-----
std = sqrt(npq) = sqrt[182*0.948] = 13.14
==========================================
Cheers,
Stan H.
===============